### v1

# ALIFE2024 template

Abstract
A major challenge when describing the origin of life is to explain how instructional information control systems emerge naturally and spontaneously from mere molecular dynamics. So far, no one has clarified how information control emerged ab initio and how primitive control mechanisms in life might have evolved, becoming increasingly refined. Based on recent experimental results showing that chemical computation does not require the presence of life-related chemistry (Dueñas-Díez and Pérez-Mercader,, 2019), we elucidate the origin and early evolution of information handling by chemical automata, from information processing (computation) to information storage (memory) and information transmission (communication). In contrast to other theories that assume the existence of initial complex structures, our narrative starts from trivial self-replicators whose interaction leads to the arising of more powerful molecular machines. By describing precisely the primordial transitions in chemistry-based computation, our metaphor is capable of explaining the above-mentioned gaps and can be translated to other models of computation, which allow us to explore biological phenomena at multiple spatial and temporal scales. At the end of our manuscript, we propose some ways to extend our ideas, including experimental validation of our theory (both in vitro and in silico).
As was explained in Walker and Davies, (2013), neither genetic-first nor metabolism-first approaches can describe completely the emergence of living organisms, as both viewpoints neglect the active and distributed nature of biological information. In this way, the real challenge when describing life’s origin is to explain how instructional information control systems emerge naturally and spontaneously from simple molecular dynamics. Thus, we explore an information-first approach to study the origins and evolution of life. Recently, it was shown that chemical reactions can be thought of as molecular recognition machines, which do not need biochemistry to mimic the behavior of automata belonging to different classes (Dueñas-Díez and Pérez-Mercader,, 2019).
The experiments carried out by Dueñas-Díez and Pérez-Mercader, (2019) (of which the second author will be one of the keynote speakers at ALIFE 2024) consisted of implementing strictly native chemical realizations of automata in their classical hierarchy. To enact a Finite-State Automaton (FSA), they used a simple bimolecular reaction in an aqueous medium. To render a one-stack Push-Down Automaton (PDA) they employed a pH-reaction network. To bring about a Turing machine in the Limited Bounded Automaton (LBA) subclass, they used the Belousov-Zhabotinsky (BZ) reaction network. As the authors have pointed out in their conclusions, these experimental results enlighten a path for the sequential/hierarchical origin of information handling in nature and the transition from non-biochemical to biochemical and extant life-related computation.
There are several theoretical frameworks which describe life as information (Chaitin,, 2012; Farnsworth et al.,, 2013; Marletto,, 2015), but none of them has explained how information handling emerged from chemical interactions. By information handling, we mean information processing (computation), information storage (memory), and information transmission (communication). In this work, we describe precisely the primordial transitions in chemistry-based computation, from the appearance of simple trivial replicators (FSAs) up to the evolution of chemical Turing machines (LBAs) and, eventually, to their combination into more powerful molecular computers.
For us, FSAs demarcate the origin of information processing (computation), PDAs define the appearance of information storage (memory) and LBAs bound the emergence of information transmission (communication). This correspondence is depicted in Fig. 1. By elucidating the origin and early evolution of information handling by chemical automata, in the next three sections we explain how information control emerged ab initio and how primitive control mechanisms in life might have evolved. We do not assume the existence of initial complex structures and our narrative can be translated to other models of computation, which in future work would allow us to explore biological phenomena at multiple spatial and temporal scales.
Figure 1: Relationship between classes of automata and the origin of the different types of information handling. The large circles represent the three forms of information handling: computation (information processing), memory (information storage), and communication (information transmission). FSAs are only capable of processing information. PDAs also have memory capacity. LBAs can additionally transmit information. For each class of automata we checklist some important properties that they exhibit given our narrative.
Information Processing
We can imagine our planet as a big parallel, quasi-universal Turing machine capable of simulating a huge variety of functions (programs). The term “quasi-universal” is due to the fact that “real” universal Turing machines are pure abstractions which need an infinite memory tape. Of course, natural resources are finite, and this is a first limitation for life to emerge. Under this perspective, the first organisms on the planet could be described as some kind of automata which were running in these primitive environments. As a result of the geophysical and geochemical constraints, which we could call the primeval ecosystem language (Pattee,, 1969), elementary bimolecular reactions emerged. These chemical reactions are capable of recognizing regular languages, a task that does not involve counting or memory (Dueñas-Díez and Pérez-Mercader,, 2019).
These trivial replicators, as Walker and Davies, (2013) call them, are simple enough to have been naturally occurring in pre-biological environments and kick off the origin of information processing. These structures do not require a memory nor a “vehicle” (in the von Neumann sense), so we can imagine the architecture of these earliest chemical machines as being conformed by a single FSA, capable of constructing itself using natural resources and a simple description. Just as $L$-systems, it is clear that the algorithmic information of these simple architectures is one that may be specified by an algorithm containing far fewer bits than the structure it describes.
In this primitive environment, it is possible to achieve self-replication (a process that differs from self-reproduction (Marletto,, 2015)) by simple bimolecular reactions (Duim and Otto,, 2017). Since it involves the unassisted formation of a covalent bond between the two reactants, the reactions will be relatively slow. Nevertheless, the autocatalytic cycle will play an increasingly dominant role (Kauffman,, 1986) if a sufficiently large amount of the template molecules is formed (Duim and Otto,, 2017). Although it has been speculated that the simple interaction between self-replicators and the environment leads to open-ended evolution, for now this is not a prerequisite.
Information Storage
So far, we have a world in which molecular FSAs coexist. These are capable of computation and self-replication but have no vehicles or memory. The diversity of molecules in the environment allows them to interact and, if possible, bind with each other. This generates systems of multiple reactions that are interconnected through common intermediate chemical species. Under this panorama, some chemical intermediates are acting as the product of one reaction and as reactant of another one. This dynamical process is essentially equivalent to the push/pop duality in one-stack PDAs (Dueñas-Díez and Pérez-Mercader,, 2019). This transition from FSAs to one-stack PDAs demarcates the appearance of analog structures, in which delocalized information storage (memory) allows for a higher evolutionary robustness in the face of changing environmental conditions (Walker and Davies,, 2013).
From a von Neumann mindset, the one-stack PDA is now some kind of primitive constructor. At this computational level, self-replication is the result of nature trying to solve a equivalent PDA-problem to the Halting problem (Sayama,, 2008). In this case, the ecosystem, a quasi-universal computer compartment, tries to predict which PDA will be the longest-running machine. This is equivalent to computing the largest number of steps taken by an $n$-state, $m$-symbol machine started on an initially non-blank tape before halting. This is of course an undecidable problem, and at the level of the PDA, when this attempts to execute the instruction of being the longest-running machine, it produces a nested loop of computation which in terms of construction translates as self-replication. This process is summarized in Fig. 2.
Figure 2: Explanation of PDA and LBA self-replication. First, the environment (described as a quasi-universal computer compartment) tries to predict which replicator will be the longest-running machine. Next, the automata attempt to execute the instruction of being the longest-running machine. This results in a nested loop of computation which in terms of construction translates as self-replication (Sayama,, 2008). Once the automata self-replicate, the cycle repeats. Undecidability explicitly leads us to have open-ended evolution. (Rocha,, 1998; Hernández-Orozco et al.,, 2018).
The undecidability behind self-replication lead us explicitly to open-ended evolution (Rocha,, 1998; Hernández-Orozco et al.,, 2018). In fact, the delocalized memory allows the PDA to store a more complex set of instructions, including the components of the chemical network conforming the PDA itself. Thus, in this evolutionary phase the system is capable of developing extra functions, which gives to the PDA the ability of self-preserving behavior and propagate mutations to its offspring. In addition, observe that the transition between FSAs and one-stack PDAs is compatible with ideas related to the appearance composomes (Segré and Lancet,, 2000; Hunding et al.,, 2006) and their maintenance (Markovitch and Lancet,, 2012).
At this stage we can talk about ante-organisms, life-like systems that are capable of self-preserving behaviour (Egbert et al.,, 2023). The minimal requirement for natural selection is that each kind of replicator produce at least one viable offspring, on average, per lifetime (Marletto,, 2015). Therefore, we can claim that so far, our ante-organisms are capable of primitive evolution (some kind of natural selection with limited heredity). Of course, it is well-known that analog-only systems are not as versatile or as robust as analog systems with digital information control (Walker and Davies,, 2013), and as such, may probably have a limited evolutionary capacity (Vasas et al.,, 2010).
Information Transmission
The next level of computational complexity are Turing machines, a class of automaton equivalent to a two or more stacks PDA (Hopcroft et al.,, 2001). Following our train of thought, there are two main, non-exclusive ways for obtaining an LBA. The first one is through the complexification of one initial PDA. Since at this stage the entities are able to propagate mutations across their offspring, it is possible that they can become more and more complex, in a way that they can develop two or more push/pop dualities. The second way is essentially through the coupling of two or more PDAs. Our model does not exclude the coexistence of multiple chemical automata with different degrees of complexity. Given an environment with finite resources promoting undecidability, these different organisms can cooperate to persist.
Because chemical automata need communication in some way to cooperate, this era launches information transmission. At this evolutionary phase, we can talk about some kind of compartmentalization, in the sense that now the chemical constructors can be made of at least two other chemical subconstructors, increasing their functionality and robustness. In this period, we might have observed the emergence of components capable of copying non-digital patterns and performing part of the manipulation of those rules. Notice that at this stage the abstract and non-physical systemic entities effectively become causal agents capable of manipulating more and more of its material substrate (Auletta et al.,, 2008). Following the mindset of (Walker and Davies,, 2013), we are having a transition from trivial to non-trivial replicators, where the latter requires an algorithm, or instruction set, of complexity comparable to the system it describes (or creates).
Since undecidability is plausible in the computational classes of PDA and LBA (Rado,, 1962), the entities can still self-replicate by the mechanism previously described (see Fig. 2). However, they are not yet capable of genetic-like Darwinian evolution, so we still have ante-organisms (Egbert et al.,, 2023). In this world of FSAs, PDAs, and LBAs interacting in a dynamical environment, the appearance of digital messengers is not totally unlikely since these structures should have had enough time to emerge as the product of some synergistic process, in much the same way that our formal rules of syntax and dictionaries were condensed from the functional usage of primitive symbols in a complex environment (Pattee,, 1969). Along these lines, digitization would have arisen as a natural outcome of the processes in reaction networks that had once been primarily analog. This fact helped automata to communicate more efficiently among them and at the same time with their offspring. Certainly, this process could have taken a long time to happen, such that digital life forms may be the only systems that survive on the long-run and are thus the only remaining product of the processes that led to life (Walker and Davies,, 2013).
The appearance of digital messengers can be seen as the implementation of descriptions that can (or cannot) interact with the primitive chemical constructors to be constructed or copied. Of course, this is not a necessary condition for life, and it will depend on which architecture we are talking about. For instance, nearly all biochemical interactions in the cell handle information in an analog format (Walker and Davies,, 2013). Now that we have a “vehicle” in the von Neumann sense, we can talk about self-reproduction (Marletto,, 2015) and proto-organisms (Egbert et al.,, 2023), entities capable of gene-based Darwinian evolution, but possibly lacking some features found in all modern organisms.
Discussion
Based on the experimental results obtained in Dueñas-Díez and Pérez-Mercader, (2019), our work explains the origin and early evolution of information handling by matter in proto-habitable environments. It is very important to mention that, although we draw inspiration from that paper, our work does not claim that the chemical reactions proposed by Dueñas-Díez and Pérez-Mercader, (2019) are the only path to life’s computation. On the contrary, our arguments imply that the origin and evolution of information handling is independent of the substrate, so that the existence of a specific chemical reaction is not necessary. The real challenge of the origin of life is not as such the existence of chemical computation, but reliable computation capable of permeating the protobiosphere. Particularly the framework proposed by Chollet, (2019) seems suitable to quantify how resilient the chemical entities proposed here are in the face of a changing environment.
Despite that our vision is consistent with the fact that symbol grounding precedes interpretation (Pattee,, 2021) and that we have used many computational concepts when describing our ideas, we did not consider symbols as fundamental entities, i.e., our approach does not presuppose an objectivist ontology. Therefore, our narrative is constructivist, and as has been shown in Stewart, (1995), it is possible to describe from life’s origin to higher-level cognition using purely a constructivist description. Here the attention of our explanation it is confined to unicellular organisms reproducing asexually in a non-biological environment, but our narrative suggests that the combination of chemical automata into nested, more powerful molecular processors would lead us eventually to the advent of organisms (Egbert et al.,, 2023), or life as we know it. In this way, we could extend our ideas to other temporal and spatial biological scales.
Since our theory is compatible with the von Neumann’s replicator-vehicle logic, it is possible under no-design laws (Marletto,, 2015), provided that the latter permit information media and enough generic resources. In this way, we have already proposed the path for the sequential/hierarchical origin of information handling in nature and the transition from non-biochemical to biochemical and extant life-related computation. Notwithstanding, even though our narrative describes how the primitive information control mechanisms in life might have evolved, we do not claim from what point life has emerged. Essentially, depending on how we define life (Benner,, 2010), we could affirm that life has started from some arbitrary point in our hypothetical description. Since we are working from a gradualist perspective of life, this is not problematic. In fact, the possibility of characterizing the different levels of complexification in life may well be related to the major transitions in evolution (Szathmáry and Smith,, 1995).
Following the last paragraph, we can arbitrarily consider that proto- or even ante-organisms (distinguished in Egbert et al., (2023)) are already alive. In fact, someone could hold that life emerged together with natural selection, or compartmentation, or agency, or any other criterium, and then we agree with Walker and Davies, (2013) when they stated that a precise point of transition from non-life to life may actually be undecidable in the logical sense. In addition, since resources are finite, it is not possible for chemistry alone to allow universal computation in the strictly formal sense (Dueñas-Díez and Pérez-Mercader,, 2019). Thus, under our paradigm, life does not need universal computation to exist. Finalizing this paragraph, at different stages of the universe, multiple places in the cosmos can serve as the quasi-universal Turing machine mentioned here. Therefore, our model is agnostic for the period life could have emerged in the universe (Loeb,, 2014; Pearce et al.,, 2018), as well with where life can occur (Abbot and Switzer,, 2011; Ballesteros et al.,, 2019).
In this fashion, with our narrative, we have solved the problem faced by Turing and von Neumann when trying to decompose and store the algorithm to build an organism in a simple sequential digital structure to be read-out by an appropriate machine. Under our portrayal, the algorithm for building an organism is not only stored in a linear digital sequence (tape), but also in the current state of the entire system. That is, we have blurred the distinction between information storage (memory) and information processing (computation) within primitive von Neumann architectures, just as life does. Even though our ideas answer the question of how information control emerged ab initio, as well as how primitive control mechanisms might evolve and become increasingly refined after an algorithmic takeover has occurred (Walker and Davies,, 2013), they can be extended in different ways and some of them are mentioned below.
The metaphor proposed here has one main advantage, which is illustrated in Fig. 3. There are well-known equivalences between different models of computation (Moore and Mertens,, 2011), which means that our ideas can be described not only in terms of sequential models (automata), but also in terms of functional and concurrent models of computation, such as $\lambda$-calculus or logic gates. The possibility of describing any algorithm by means of a Turing machine (Church-Turing thesis) together with the equivalences between the different models of computation, makes it feasible that our ideas allow us to portray life not only in a primordial way but at other evolutionary stages. For instance, here we did not explore the thermodynamic aspects of life, but if living systems are qualitatively (and quantitatively) well described by the metaphor here presented, we can use different models of computation to describe the thermodynamics of different facets of life (Zenil et al.,, 2012; Farnsworth et al.,, 2013; Kolchinsky and Wolpert,, 2020; Wolpert and Kolchinsky,, 2020). Similarly, since it is affordable to characterize von Neumman’s replicator-vehicle logic at the quantum level (Marletto,, 2015), our model could explore the microscopic aspects in the origin and evolution of life (Abbott et al.,, 2008).
Figure 3: Main advantage of our narrative outlined. If any real-world computation can be translated into an equivalent computation involving a Turing machine, we can use different models of computation to describe different biological phenomena. For example, we could explain the molecular processes of the cell using some functional model such as lambda calculus or general recursive functions. Similarly, we can resort to recurrent models such as cellular automata or logic gates to describe the interaction between cells. At the holobiont level we can use sequential models (automata, like the ones described here) to characterize the behavior of more complex organisms. This would allow us to study life at different nested, spatiotemporal layers and is in agreement with the idea that there is no privileged level of causation (Noble,, 2012).
In order to describe life as information at other spatial and temporal scales, we can integrate our narrative in other theses such as Farnsworth et al., (2013), where persistence, context, and effective information play an important role to describe life at all levels from a unifying outlook. Again, given that under our perspective life is describable at all extents, there should not be privileged level of causation (Noble,, 2012). This opens the possibility to characterize phenomena like downward causation (Flack,, 2017) from our theoretical framework, and even if the informational-causal narrative is accepted as a mere façon de parler, our metaphor is still compatible with theories where only material objects enjoy true causal efficacy (Walker et al.,, 2012).
Future work
For the purpose of testing our hypotheses and not to leave them as a mere speculative model (Pattee,, 2021), we can appeal to experiments in vitro and in silico. Of the many approaches that exist in synthetic biology (Hanczyc,, 2020), protocells offers a comprehensive resource on current attempts to create simple forms of life from scratch in the laboratory (Deamer,, 2008). In fact, using the BZ reaction (equivalent to a LBA) it is possible to devise active synthetic polymer vesicles, which self-reproduce (Bastakoti and Perez-Mercader,, 2017) and evolve (Katla et al.,, 2023). Is it possible to synthesize vesicles using a pH chemical reaction (equivalent to a PDA) or a precipitation reaction (equivalent to a FSA)? If yes, what are the abilities of these new vesicles? If not, what type of structures can we synthesize and what are their abilities? The same research group has obtained three reaction-diffusion equations that are capable of representing information processing, metabolism, self-replication and evolution (Muñuzuri and Pérez-Mercader,, 2022). The solutions to these equations show the phenomenology of a generic living system. To obtain these equations the authors implemented each of the four properties (information processing, metabolism, self-replication and evolution) one by one. We can eliminate any of these assumptions, modifying the equations. For example, we can start by a set of equations which only allow information processing and self-replication. What kind of phenomenology is obtained by solving this new system of equations? Are these results consistent with our theory? The answer to these questions can tell us how valid are our ideas.
Recently, a new relational biology model which looks for the minimal prototype for organized systems has been developed (Márquez-Zacarías,, 2024). This theoretical framework allows to generate a taxonomy for the different types of organization observed, and in comparison to other theories, it does not presuppose the existence of a boundary between the environment and the organism, nor the extant of complex building blocks ab initio. To perform simulations in this framework, we need as input a set of entities and relations. We can use as input the chemical components and reactions proposed in (Dueñas-Díez and Pérez-Mercader,, 2019) for every class of chemical automata and observe which types of minimal paths of organization are formed. This is something similar to the methodology proposed by Fontana and Buss, (1994), which was refined later in Szathmáry, (1995). Are these organizations able to self-reproduce, compute, have memory, or communicate? Again, the answer to this question can tell us how valid are our ideas.
Finally, as we have observed, there is a parallelism between the classes of chemical automata in (Dueñas-Díez and Pérez-Mercader,, 2019) and the classification of entities characterized in (Egbert et al.,, 2023). Table 1 summarizes this correspondence. At every stage, our description of how the evolutionary units evolve allows them to select and/or modifying their environment using some form of basic replication. The viability-based behaviour proposed in (Egbert et al.,, 2023) increases the feasibility, diversity, and robustness of biological entities, allowing for a wider variety of organismic forms capable of persisting in a wider variety of environments, and it is also easy for nature to implement. Together with our description of how information handling appeared, viability-based behaviour it is a good candidate to explore how agency has emerged and evolved. In this way, instead of creating new theoretical frameworks and computational tools to study phenomena such as agency, life, intelligence, identity, and cognition, we promote the unification of what already exists by looking for global equivalences among different formalisms.
Types of organisms Corresponding automata
Ante-organisms PDA
Proto-organisms LBA with digital messengers
Organisms More complex nested automata
Table 1: Resemblance between the classification of organisms proposed by Egbert et al., (2023) and their respective characterization under our framework in terms of automata.

### v2

@@
 Abstract
-A major challenge when describing the origin of life is to explain how instructional information control systems emerge naturally and spontaneously from mere molecular dynamics.
+A major challenge when describing the origin of life is to explain “how instructional information control systems emerge naturally and spontaneously from mere molecular dynamics” (Walker and Davies,, 2013).
 So far, no one has clarified how information control emerged ab initio and how primitive control mechanisms in life might have evolved, becoming increasingly refined.
 Based on recent experimental results showing that chemical computation does not require the presence of life-related chemistry (Dueñas-Díez and Pérez-Mercader,, 2019), we elucidate the origin and early evolution of information handling by chemical automata, from information processing (computation) to information storage (memory) and information transmission (communication).
-In contrast to other theories that assume the existence of initial complex structures, our narrative starts from trivial self-replicators whose interaction leads to the arising of more powerful molecular machines.
-By describing precisely the primordial transitions in chemistry-based computation, our metaphor is capable of explaining the above-mentioned gaps and can be translated to other models of computation, which allow us to explore biological phenomena at multiple spatial and temporal scales.
+In contrast to other theories that assume the existence of initial complex structures, our representation starts from trivial self-replicators whose interaction leads to the arising of more powerful molecular machines.
+By describing precisely the primordial transitions in chemistry-based computation, our framework is capable of explaining the above-mentioned gaps and can be translated to other models of computation, which allow us to explore biological phenomena at multiple spatial and temporal scales.
 At the end of our manuscript, we propose some ways to extend our ideas, including experimental validation of our theory (both in vitro and in silico).
-As was explained in Walker and Davies, (2013), neither genetic-first nor metabolism-first approaches can describe completely the emergence of living organisms, as both viewpoints neglect the active and distributed nature of biological information.
-In this way, the real challenge when describing life’s origin is to explain how instructional information control systems emerge naturally and spontaneously from simple molecular dynamics.
+Introduction
+Leaving aside the philosophical question about the origin of information or the chicken-and-egg problem between information and matter/energy, it is possible to have an even deeper understanding of phenomena such as life or cognition if we assume that information is anything that an entity can sense, perceive, or observe.
+(Gershenson,, 2012).
+Thus, following the line of thought proposed by Pattee, (1969), we can establish the primordial ecosystem as the first information-containing architecture in our planet, whose primitive geochemical cycles gave way to a well-defined syntax—based on chemistry— spontaneously without the need for genetic instructions or initial metabolic control.
+This gave way to molecular Darwinism, a phenomenon that has been well documented for several decades (Küppers,, 1990) and whose formalization was refined the last decade by Marletto, (2015).
+As was explained in Walker and Davies, (2013), neither genetic-first nor metabolism-first approaches can describe completely the emergence of living organisms, as both viewpoints “neglect the active (algorithmic or instructional) and distributed nature of biological information” (Walker and Davies,, 2013).
+In this way, “the real challenge of life’s origin is thus to explain how instructional information control systems emerge naturally and spontaneously from mere molecular dynamics” (Walker and Davies,, 2013).
 Thus, we explore an information-first approach to study the origins and evolution of life.
-Recently, it was shown that chemical reactions can be thought of as molecular recognition machines, which do not need biochemistry to mimic the behavior of automata belonging to different classes (Dueñas-Díez and Pérez-Mercader,, 2019).
-The experiments carried out by Dueñas-Díez and Pérez-Mercader, (2019) (of which the second author will be one of the keynote speakers at ALIFE 2024) consisted of implementing strictly native chemical realizations of automata in their classical hierarchy.
+Recently, it was shown that chemical reactions can be thought of as molecular recognition machines, which do not need biochemistry to mimic the behavior of recognizing automata belonging to different classes in Chomsky’s hierarchy (Dueñas-Díez and Pérez-Mercader,, 2019).
+The experiments carried out by Dueñas-Díez and Pérez-Mercader, (2019) consisted of implementing strictly “one-pot-reactor” chemical realizations of automata for the recognition of well-known languages at different levels in the Chomsky hierarchy.
 To enact a Finite-State Automaton (FSA), they used a simple bimolecular reaction in an aqueous medium.
-To render a one-stack Push-Down Automaton (PDA) they employed a pH-reaction network.
-To bring about a Turing machine in the Limited Bounded Automaton (LBA) subclass, they used the Belousov-Zhabotinsky (BZ) reaction network.
-As the authors have pointed out in their conclusions, these experimental results enlighten a path for the sequential/hierarchical origin of information handling in nature and the transition from non-biochemical to biochemical and extant life-related computation.
+Such reaction was capable of recognizing a regular language.
+To render a one-stack Push-Down Automaton (PDA) they employed a pH-reaction network, being capable of recognizing a context-free language.
+To bring about a Turing machine in the Linear Bounded Automaton (LBA) subclass, they used the Belousov-Zhabotinsky (BZ) reaction network.
+In this case, the BZ reaction was able to successfully recognize a context-sensitive language.
+As the authors have pointed out in their conclusions, these experimental results enlighten “a path for the sequential/hierarchical origin of information handling in nature and the transition from non-biochemical to biochemical and extant life-related computation” (Dueñas-Díez and Pérez-Mercader,, 2019).
 There are several theoretical frameworks which describe life as information (Chaitin,, 2012; Farnsworth et al.,, 2013; Marletto,, 2015), but none of them has explained how information handling emerged from chemical interactions.
 By information handling, we mean information processing (computation), information storage (memory), and information transmission (communication).
-In this work, we describe precisely the primordial transitions in chemistry-based computation, from the appearance of simple trivial replicators (FSAs) up to the evolution of chemical Turing machines (LBAs) and, eventually, to their combination into more powerful molecular computers.
-For us, FSAs demarcate the origin of information processing (computation), PDAs define the appearance of information storage (memory) and LBAs bound the emergence of information transmission (communication).
+As will be discussed below, such information handling also includes the syntactic, semantic and pragmatic flavors of information.
+In this work, we describe precisely the primordial transitions in chemistry-based computation, from the appearance of trivial replicators (simple bimolecular reactions, equivalent to FSAs) in the environment up to the evolution of approximated chemical Turing machines (complex chemical networks, represented here by LBAs) and, eventually, to their combination into more powerful molecular computers, capable of performing Bayesian inference (Poole et al.,, 2017) and associative learning (Bartlett and Louapre,, 2022), which could have led to the development of autonomy (Kirchhoff et al.,, 2018) and high-level cognition (Levin,, 2019).
+For us, the presence of FSAs in the environment demarcates the origin of information processing (computation), PDAs define the appearance of information storage (memory) by the environment and LBAs bound the emergence of information transmission (communication) through different entities in the environment.
 This correspondence is depicted in Fig.
@@
 The large circles represent the three forms of information handling: computation (information processing), memory (information storage), and communication (information transmission).
-FSAs are only capable of processing information.
-PDAs also have memory capacity.
-LBAs can additionally transmit information.
-For each class of automata we checklist some important properties that they exhibit given our narrative.
+FSAs allow the environment to process information.
+PDAs permit a primitive form of memory capacity by the environment.
+LBAs empower communication between different entities in the environment.
+For each class of automata we checklist some important properties that they grant to the ecosystem given our framework.
 Information Processing
-We can imagine our planet as a big parallel, quasi-universal Turing machine capable of simulating a huge variety of functions (programs).
-The term “quasi-universal” is due to the fact that “real” universal Turing machines are pure abstractions which need an infinite memory tape.
-Of course, natural resources are finite, and this is a first limitation for life to emerge.
-Under this perspective, the first organisms on the planet could be described as some kind of automata which were running in these primitive environments.
-As a result of the geophysical and geochemical constraints, which we could call the primeval ecosystem language (Pattee,, 1969), elementary bimolecular reactions emerged.
+As has been suggested by Chaitin, (2012), think of our planet as a big parallel, quasi-universal Turing machine capable of simulating a huge variety of functions (programs) can give us novel insights about evolution.
+Here, the term “quasi-universal” is due to the fact that “real” universal Turing machines are pure abstractions which need an infinite memory tape.
+Natural resources are finite, and this is a first limitation for life to emerge.
+Extrapolating the ideas suggested in Hernández-Orozco et al., (2018) and inspired by the evolutionary model proposed in von Neumann, (1966), the first organisms on the planet could be described as some kind of primitive automata which were assembled and executed by these primeval habitats.
+As noted above, the primordial geophysical and geochemical constraints generated the primeval ecosystem language (Pattee,, 1969), which was interpreted by such primitive entities.
+The cyclic nature of such processes allows for a robust flow of information—abolishing the fragility or deterioration of it—, resulting in the emergence of elementary autocatalytic bimolecular reactions.
 These chemical reactions are capable of recognizing regular languages, a task that does not involve counting or memory (Dueñas-Díez and Pérez-Mercader,, 2019).
+By the good regulator theorem (Conant and Ross Ashby,, 1970), some of the elementary bimolecular reactions in that primeval environment must have a cyclic nature similar to that of geologic phenomena.
+Thus, it is possible to achieve self-replication by simple bimolecular reactions.
+Since it involves the unassisted formation of a covalent bond between the two reactants, the reactions will be relatively slow (Duim and Otto,, 2017).
+Nevertheless, the autocatalytic cycle will play an increasingly dominant role (Kauffman,, 1986) if a sufficiently large amount of the template molecules is formed (Duim and Otto,, 2017), which in part could also promote self-replication.
+Self-replication differs from self-reproduction in the existence of a copier $C$ acting on raw material (Szathmáry,, 1995).
+As has been well argued by Marletto, (2015): “for looser replicators, such as crystals, or short RNA strands and autocatalytic cycles involved in the origin of life, the copier $C$ is null, i.e.
+implicit in the laws of physics”.
+Although it has been speculated that the simple interaction between self-replicators and the environment leads to open-ended evolution, for us this is not a prerequisite.
 These trivial replicators, as Walker and Davies, (2013) call them, are simple enough to have been naturally occurring in pre-biological environments and kick off the origin of information processing.
-These structures do not require a memory nor a “vehicle” (in the von Neumann sense), so we can imagine the architecture of these earliest chemical machines as being conformed by a single FSA, capable of constructing itself using natural resources and a simple description.
+These structures do not require a memory nor a “vehicle”—in the von Neumann sense (von Neumann,, 1966)—, so we can imagine the architecture of these earliest chemical machines as being represented by autocatalytic bimolecular chemical reactions, which under our framework are single FSAs (Dueñas-Díez and Pérez-Mercader,, 2019), capable of constructing themselves using natural resources and a simple description encoded in the environment.
 Just as $L$-systems, it is clear that the algorithmic information of these simple architectures is one that may be specified by an algorithm containing far fewer bits than the structure it describes.
-In this primitive environment, it is possible to achieve self-replication (a process that differs from self-reproduction (Marletto,, 2015)) by simple bimolecular reactions (Duim and Otto,, 2017).
-Since it involves the unassisted formation of a covalent bond between the two reactants, the reactions will be relatively slow.
-Nevertheless, the autocatalytic cycle will play an increasingly dominant role (Kauffman,, 1986) if a sufficiently large amount of the template molecules is formed (Duim and Otto,, 2017).
-Although it has been speculated that the simple interaction between self-replicators and the environment leads to open-ended evolution, for now this is not a prerequisite.
+At this point, it is important to note that the different dimensions of the concept of information (syntactic, semantic and pragmatic) cannot be separated from each other (Küppers,, 1990).
+Although the existence of semantic (or pragmatic) information in such a primordial state of life may seem far-fetched, it must first be remembered that meaning is a multi-scale phenomenon (Fields and Levin,, 2020).
+According to Küppers, (1990): “the protosemantics of biological information is clearly defined by the ability of the biological information carrier to reproduce as quickly as possible while maintaining high accuracy and stability”, which is fully in line with our hypothesis.
+In fact, it has been shown that it is possible to approximate Bayesian inference using simple bimolecular reactions (Poole et al.,, 2017), which could have promoted the decoding of symbol-to-meaning information and subsequently developed semantic closure (Pattee,, 1995).
 Information Storage
-So far, we have a world in which molecular FSAs coexist.
+So far, we have a world in which bimolecular reactions (represented here by FSAs) coexist.
 These are capable of computation and self-replication but have no vehicles or memory.
@@
 This generates systems of multiple reactions that are interconnected through common intermediate chemical species.
-Under this panorama, some chemical intermediates are acting as the product of one reaction and as reactant of another one.
-This dynamical process is essentially equivalent to the push/pop duality in one-stack PDAs (Dueñas-Díez and Pérez-Mercader,, 2019).
-This transition from FSAs to one-stack PDAs demarcates the appearance of analog structures, in which delocalized information storage (memory) allows for a higher evolutionary robustness in the face of changing environmental conditions (Walker and Davies,, 2013).
-From a von Neumann mindset, the one-stack PDA is now some kind of primitive constructor.
-At this computational level, self-replication is the result of nature trying to solve a equivalent PDA-problem to the Halting problem (Sayama,, 2008).
-In this case, the ecosystem, a quasi-universal computer compartment, tries to predict which PDA will be the longest-running machine.
+Under this panorama, some chemical intermediates are acting as the product of one reaction and as reactant of another one, forming networks of even more complex chemical reactions.
+This dynamical process is identical to the premise of autocatalysis and is essentially equivalent to the push/pop duality in one-stack PDAs described in (Dueñas-Díez and Pérez-Mercader,, 2019).
+The transition from FSAs (simple bimolecular reactions) to one-stack PDAs (networks of chemical reactions) demarcates the appearance of a primitive information storage (memory) by the environment, because it allows for the recognition of more complex sequences of information (Hopcroft et al.,, 2001) and thus it empowers a higher evolutionary robustness of such information in the face of changing environmental conditions (Walker and Davies,, 2013).
+From a von Neumann mindset, the one-stack PDA is now some kind of primitive constructor, because we have a chemical network that might be able to reproduce itself including the components that conform it.
+At this computational level, self-replication can be seen as the result of nature trying to solve a equivalent PDA-problem to the Halting problem (Sayama,, 2008).
+Following our computational train of thought, the ecosystem, a quasi-universal computer compartment, “is constantly approximating solutions to the adaptation problem by means of survivability” (Hernández-Orozco et al.,, 2018).
+In this way, nature pursuits of the busy beaver values (longest-running automata) in the environment, which is identical to the search for the longest-surviving organisms.
+At this particular evolutionary stage, nature is looking for which PDA will be the longest running machine, which is interchangeable to find the most robust primitive information storage for the environment.
 This is equivalent to computing the largest number of steps taken by an $n$-state, $m$-symbol machine started on an initially non-blank tape before halting.
-This is of course an undecidable problem, and at the level of the PDA, when this attempts to execute the instruction of being the longest-running machine, it produces a nested loop of computation which in terms of construction translates as self-replication.
+This is of course an undecidable problem comparable to a busy beaver game (Rado,, 1962), and at the level of the PDA, when this attempts to execute the instruction of being the longest-running machine, it produces a nested loop of computation which in terms of construction translates as self-replication (Sayama,, 2008).
 This process is summarized in Fig.
@@
 Figure 2: Explanation of PDA and LBA self-replication.
-First, the environment (described as a quasi-universal computer compartment) tries to predict which replicator will be the longest-running machine.
-Next, the automata attempt to execute the instruction of being the longest-running machine.
+First, the environment (described as a quasi-universal computer compartment) approximates solutions to the adaptation problem by means of survivability, which guides the search for the longest-surviving organisms.
+Next, the automata (a chemical network) attempt to execute the instruction of being the longest-running machine.
 This results in a nested loop of computation which in terms of construction translates as self-replication (Sayama,, 2008).
@@
 Undecidability explicitly leads us to have open-ended evolution.
-(Rocha,, 1998; Hernández-Orozco et al.,, 2018).
-The undecidability behind self-replication lead us explicitly to open-ended evolution (Rocha,, 1998; Hernández-Orozco et al.,, 2018).
-In fact, the delocalized memory allows the PDA to store a more complex set of instructions, including the components of the chemical network conforming the PDA itself.
+(Hernández-Orozco et al.,, 2018).
+The undecidability behind self-replication lead us explicitly to open-ended evolution as a result of evolution (Pattee and Sayama,, 2019), and as Hernández-Orozco et al., (2018) has pointed out, systems that exhibit (strong) open-ended evolution must be undecidable.
+The PDA allows the ecosystem to store a more complex set of instructions—including the components of the chemical network that makes up the PDA itself—in a delocalized way.
 Thus, in this evolutionary phase the system is capable of developing extra functions, which gives to the PDA the ability of self-preserving behavior and propagate mutations to its offspring.
+This is consistent with the fact that a simple catalytic reaction network may be able of encoding correlations in environmental variables, evolving some kind of associative learning (Bartlett and Louapre,, 2022).
+Following the law of requisite variety (Ashby,, 1956), these chemical networks evolve concurrently to the environment in which they were developed, producing more semantic information which translates as a primitive degree of agency (Kolchinsky and Wolpert,, 2018).
 In addition, observe that the transition between FSAs and one-stack PDAs is compatible with ideas related to the appearance composomes (Segré and Lancet,, 2000; Hunding et al.,, 2006) and their maintenance (Markovitch and Lancet,, 2012).
 At this stage we can talk about ante-organisms, life-like systems that are capable of self-preserving behaviour (Egbert et al.,, 2023).
-The minimal requirement for natural selection is that each kind of replicator produce at least one viable offspring, on average, per lifetime (Marletto,, 2015).
-Therefore, we can claim that so far, our ante-organisms are capable of primitive evolution (some kind of natural selection with limited heredity).
-Of course, it is well-known that analog-only systems are not as versatile or as robust as analog systems with digital information control (Walker and Davies,, 2013), and as such, may probably have a limited evolutionary capacity (Vasas et al.,, 2010).
+Up to this point we have a resource-limited ecosystem inhabited by chemical networks that compete, self-replicate and mutate (through small variations in their composition) in such a way that each kind of replicator produce at least one viable offspring, on average, per lifetime.
+Therefore, we can claim that so far, our ante-organisms are capable of primitive evolution (Marletto,, 2015), which is a kind of natural selection with limited inheritance caused by low robustness and accuracy in replicating the information stored in the chemical network by the environment.
+However, according to Walker and Davies, (2013): “analogue-only systems are not as versatile or as robust as analogue systems with digital information control, and as such, may probably have a limited evolutionary capacity (Vasas et al.,, 2010)”.
 Information Transmission
@@
 The first one is through the complexification of one initial PDA.
-Since at this stage the entities are able to propagate mutations across their offspring, it is possible that they can become more and more complex, in a way that they can develop two or more push/pop dualities.
+Since at this stage the entities are able to propagate mutations across their offspring (represented by variations in the structures of the chemical networks), it is possible that they can become more and more complex, in a way that they can develop two or more push/pop dualities.
 The second way is essentially through the coupling of two or more PDAs.
 Our model does not exclude the coexistence of multiple chemical automata with different degrees of complexity.
-Given an environment with finite resources promoting undecidability, these different organisms can cooperate to persist.
+Given an environment with finite resources promoting survivability, these different organisms can cooperate to persist.
+In both cases we obtain even more complex chemical networks that allow the identification of more abstract sequences of information, resulting in an even more sophisticated information storage system on the part of the environment.
 Because chemical automata need communication in some way to cooperate, this era launches information transmission.
-At this evolutionary phase, we can talk about some kind of compartmentalization, in the sense that now the chemical constructors can be made of at least two other chemical subconstructors, increasing their functionality and robustness.
-In this period, we might have observed the emergence of components capable of copying non-digital patterns and performing part of the manipulation of those rules.
+Given a set of selfish agents minimizing their stress (surprise) and competing for information, it is possible to develop cooperative homeostatic loops which improve the predictive capabilities of the entities and gives rise to functional relationships (Levin,, 2019; Heins et al.,, 2024).
+Such infotaxis may have resulted in some kind of compartmentalization, in the sense that now the chemical constructors can be made of at least two chemical networks, increasing their functionality and robustness.
 Notice that at this stage the abstract and non-physical systemic entities effectively become causal agents capable of manipulating more and more of its material substrate (Auletta et al.,, 2008).
 Following the mindset of (Walker and Davies,, 2013), we are having a transition from trivial to non-trivial replicators, where the latter requires an algorithm, or instruction set, of complexity comparable to the system it describes (or creates).
+In the same way, we can imagine this metamorphosis from reactive to representational systems (Zenil et al.,, 2012).
 Since undecidability is plausible in the computational classes of PDA and LBA (Rado,, 1962), the entities can still self-replicate by the mechanism previously described (see Fig.
@@
 This fact helped automata to communicate more efficiently among them and at the same time with their offspring.
-Certainly, this process could have taken a long time to happen, such that digital life forms may be the only systems that survive on the long-run and are thus the only remaining product of the processes that led to life (Walker and Davies,, 2013).
+Certainly, this process could have taken a long time to happen, such that digital life forms may be “the only systems that survive in the long-run and are thus the only remaining product of the processes that led to life” (Walker and Davies,, 2013).
 The appearance of digital messengers can be seen as the implementation of descriptions that can (or cannot) interact with the primitive chemical constructors to be constructed or copied.
@@
 Discussion
-Based on the experimental results obtained in Dueñas-Díez and Pérez-Mercader, (2019), our work explains the origin and early evolution of information handling by matter in proto-habitable environments.
+Based on the experimental results obtained in Dueñas-Díez and Pérez-Mercader, (2019), our theory explains the origin and early evolution of information handling by matter in proto-habitable environments.
 It is very important to mention that, although we draw inspiration from that paper, our work does not claim that the chemical reactions proposed by Dueñas-Díez and Pérez-Mercader, (2019) are the only path to life’s computation.
 On the contrary, our arguments imply that the origin and evolution of information handling is independent of the substrate, so that the existence of a specific chemical reaction is not necessary.
-The real challenge of the origin of life is not as such the existence of chemical computation, but reliable computation capable of permeating the protobiosphere.
-Particularly the framework proposed by Chollet, (2019) seems suitable to quantify how resilient the chemical entities proposed here are in the face of a changing environment.
-Despite that our vision is consistent with the fact that symbol grounding precedes interpretation (Pattee,, 2021) and that we have used many computational concepts when describing our ideas, we did not consider symbols as fundamental entities, i.e., our approach does not presuppose an objectivist ontology.
-Therefore, our narrative is constructivist, and as has been shown in Stewart, (1995), it is possible to describe from life’s origin to higher-level cognition using purely a constructivist description.
-Here the attention of our explanation it is confined to unicellular organisms reproducing asexually in a non-biological environment, but our narrative suggests that the combination of chemical automata into nested, more powerful molecular processors would lead us eventually to the advent of organisms (Egbert et al.,, 2023), or life as we know it.
-In this way, we could extend our ideas to other temporal and spatial biological scales.
-Since our theory is compatible with the von Neumann’s replicator-vehicle logic, it is possible under no-design laws (Marletto,, 2015), provided that the latter permit information media and enough generic resources.
+The real challenge of the origin of life is not the existence of such chemical computation, but reliable computation capable of permeating the protobiosphere.
+Particularly, the framework proposed by Chollet, (2019) seems suitable to quantify how resilient the chemical entities proposed here are in the face of a changing environment.
+In order for there to be such automata capable of interpreting and storing the information contained in the environment, there must first be a primordial language (conformed by chemicals) in the environment that allows the realization of such automata.
+Therefore, our vision is consistent with the fact that symbol grounding precedes interpretation (Pattee,, 2021).
+We have used many computational concepts when describing our ideas, however, our theory points to the fact that life and cognition are two intertwined, parallel phenomena.
+As has been shown in Stewart, (1995), it is possible to describe from life’s origin to higher-level cognition using purely a constructivist description.
+Being compatible with the free energy principle, in this paper we have developed a computational enactivist (Korbak,, 2021) theoretical framework capable of being as explanatory as a strictly constructivist approach.
+To justify this last assertion, let us take into consideration that the attention of our explanation it is confined to unicellular organisms reproducing asexually in a non-biological environment, but our framework suggests that the combination of chemical automata into nested, more powerful molecular processors would lead us eventually to the advent of organisms (Egbert et al.,, 2023), or “Life as we know it” (Friston,, 2013).
+In this way, we could extend our ideas to other temporal and spatial biological scales, where molecular machines are exponentially more complex, allowing a Bayesian inference (Poole et al.,, 2017) that would later reach different degrees of autonomy as an evolutionary advantage (Kirchhoff et al.,, 2018).
+Additionally, we could describe and analyze the evolution of open-endedness (Pattee and Sayama,, 2019).
+Since our theory is compatible with the von Neumann’s replicator-vehicle logic (von Neumann,, 1966), it is possible under no-design laws (Marletto,, 2015), provided that the latter permit information media and enough generic resources.
 In this way, we have already proposed the path for the sequential/hierarchical origin of information handling in nature and the transition from non-biochemical to biochemical and extant life-related computation.
@@
 Essentially, depending on how we define life (Benner,, 2010), we could affirm that life has started from some arbitrary point in our hypothetical description.
-Since we are working from a gradualist perspective of life, this is not problematic.
-In fact, the possibility of characterizing the different levels of complexification in life may well be related to the major transitions in evolution (Szathmáry and Smith,, 1995).
-Following the last paragraph, we can arbitrarily consider that proto- or even ante-organisms (distinguished in Egbert et al., (2023)) are already alive.
+Since we are working from a gradualist perspective of life, this is not problematic (Küppers,, 1990).
+In fact, the possibility of characterizing the different levels of complexification in life may well be related to the major transitions in evolution (Szathmáry and Smith,, 1995; West et al.,, 2015).
+Following the last paragraph, we can arbitrarily consider that proto- or even ante-organisms—distinguished in Egbert et al., (2023)—are already alive.
 In fact, someone could hold that life emerged together with natural selection, or compartmentation, or agency, or any other criterium, and then we agree with Walker and Davies, (2013) when they stated that a precise point of transition from non-life to life may actually be undecidable in the logical sense.
-In addition, since resources are finite, it is not possible for chemistry alone to allow universal computation in the strictly formal sense (Dueñas-Díez and Pérez-Mercader,, 2019).
-Thus, under our paradigm, life does not need universal computation to exist.
 Finalizing this paragraph, at different stages of the universe, multiple places in the cosmos can serve as the quasi-universal Turing machine mentioned here.
 Therefore, our model is agnostic for the period life could have emerged in the universe (Loeb,, 2014; Pearce et al.,, 2018), as well with where life can occur (Abbot and Switzer,, 2011; Ballesteros et al.,, 2019).
-In this fashion, with our narrative, we have solved the problem faced by Turing and von Neumann when trying to decompose and store the algorithm to build an organism in a simple sequential digital structure to be read-out by an appropriate machine.
-Under our portrayal, the algorithm for building an organism is not only stored in a linear digital sequence (tape), but also in the current state of the entire system.
+In this fashion, with our model, we have solved the problem faced by Turing when trying to decompose and store the algorithm to build an organism in a simple sequential digital structure to be read-out by an appropriate machine (Brenner,, 2012).
+Under our portrayal, the algorithm for building an organism is not only stored in a linear digital sequence (tape), but also in the current state of the entire system, which is clearly inspired by von Neumann, (1966).
 That is, we have blurred the distinction between information storage (memory) and information processing (computation) within primitive von Neumann architectures, just as life does.
 Even though our ideas answer the question of how information control emerged ab initio, as well as how primitive control mechanisms might evolve and become increasingly refined after an algorithmic takeover has occurred (Walker and Davies,, 2013), they can be extended in different ways and some of them are mentioned below.
-The metaphor proposed here has one main advantage, which is illustrated in Fig.
+The main advantage of our theory is illustrated in Fig.
 3.
-There are well-known equivalences between different models of computation (Moore and Mertens,, 2011), which means that our ideas can be described not only in terms of sequential models (automata), but also in terms of functional and concurrent models of computation, such as $\lambda$-calculus or logic gates.
+There are well-known equivalences between different models of computation (Moore and Mertens,, 2011), which means that our approach can be described not only in terms of sequential models (automata), but also in terms of functional and concurrent models of computation, such as $\lambda$-calculus or logic gates.
 The possibility of describing any algorithm by means of a Turing machine (Church-Turing thesis) together with the equivalences between the different models of computation, makes it feasible that our ideas allow us to portray life not only in a primordial way but at other evolutionary stages.
-For instance, here we did not explore the thermodynamic aspects of life, but if living systems are qualitatively (and quantitatively) well described by the metaphor here presented, we can use different models of computation to describe the thermodynamics of different facets of life (Zenil et al.,, 2012; Farnsworth et al.,, 2013; Kolchinsky and Wolpert,, 2020; Wolpert and Kolchinsky,, 2020).
+For instance, here we did not explore the thermodynamic aspects of life, but if living systems are qualitatively (and quantitatively) well described by the framework here presented, we can use different models of computation to describe the thermodynamics of different facets of life (Zenil et al.,, 2012; Farnsworth et al.,, 2013; Kolchinsky and Wolpert,, 2020; Wolpert and Kolchinsky,, 2020).
 Similarly, since it is affordable to characterize von Neumman’s replicator-vehicle logic at the quantum level (Marletto,, 2015), our model could explore the microscopic aspects in the origin and evolution of life (Abbott et al.,, 2008).
@@
 In order to describe life as information at other spatial and temporal scales, we can integrate our narrative in other theses such as Farnsworth et al., (2013), where persistence, context, and effective information play an important role to describe life at all levels from a unifying outlook.
-Again, given that under our perspective life is describable at all extents, there should not be privileged level of causation (Noble,, 2012).
+Given that under our perspective life is describable at all extents, there should not be privileged level of causation (Noble,, 2012).
 This opens the possibility to characterize phenomena like downward causation (Flack,, 2017) from our theoretical framework, and even if the informational-causal narrative is accepted as a mere façon de parler, our metaphor is still compatible with theories where only material objects enjoy true causal efficacy (Walker et al.,, 2012).
 Future work
-For the purpose of testing our hypotheses and not to leave them as a mere speculative model (Pattee,, 2021), we can appeal to experiments in vitro and in silico.
+For the purpose of testing our hypotheses and not to leave them as an speculative model (Pattee,, 2021), we can appeal to experiments in vitro and in silico.
 Of the many approaches that exist in synthetic biology (Hanczyc,, 2020), protocells offers a comprehensive resource on current attempts to create simple forms of life from scratch in the laboratory (Deamer,, 2008).
@@
 To perform simulations in this framework, we need as input a set of entities and relations.
-We can use as input the chemical components and reactions proposed in (Dueñas-Díez and Pérez-Mercader,, 2019) for every class of chemical automata and observe which types of minimal paths of organization are formed.
-This is something similar to the methodology proposed by Fontana and Buss, (1994), which was refined later in Szathmáry, (1995).
+We can use as input the chemical components and reactions proposed in Dueñas-Díez and Pérez-Mercader, (2019) for every class of chemical automata and observe which types of minimal paths of organization are formed.
+This is something similar to the methodology proposed by Fontana and Buss, (1994), which was refined later by Szathmáry, (1995).
 Are these organizations able to self-reproduce, compute, have memory, or communicate?
 Again, the answer to this question can tell us how valid are our ideas.
-Finally, as we have observed, there is a parallelism between the classes of chemical automata in (Dueñas-Díez and Pérez-Mercader,, 2019) and the classification of entities characterized in (Egbert et al.,, 2023).
+Finally, as we have observed, there is a parallelism between the classes of chemical automata in Dueñas-Díez and Pérez-Mercader, (2019) and the classification of entities characterized in Egbert et al., (2023).
 Table 1 summarizes this correspondence.
 At every stage, our description of how the evolutionary units evolve allows them to select and/or modifying their environment using some form of basic replication.
-The viability-based behaviour proposed in (Egbert et al.,, 2023) increases the feasibility, diversity, and robustness of biological entities, allowing for a wider variety of organismic forms capable of persisting in a wider variety of environments, and it is also easy for nature to implement.
-Together with our description of how information handling appeared, viability-based behaviour it is a good candidate to explore how agency has emerged and evolved.
+The viability-based behaviour proposed in Egbert et al., (2023) increases the feasibility, diversity, and robustness of biological entities, “allowing for a wider variety of organismic forms capable of persisting in a wider variety of environments” (Egbert et al.,, 2023), and it is also easy for nature to implement.
+Together with our description of how information handling appeared, viability-based behaviour it is a good candidate to explore how agency has emerged and evolved, or to model scale-free cognition (Levin,, 2019).
 In this way, instead of creating new theoretical frameworks and computational tools to study phenomena such as agency, life, intelligence, identity, and cognition, we promote the unification of what already exists by looking for global equivalences among different formalisms.
@@
 Table 1: Resemblance between the classification of organisms proposed by Egbert et al., (2023) and their respective characterization under our framework in terms of automata.
+Acknowledgements
+The authors acknowledge Luis M. Rocha and three anonymous reviewers for their valuable comments during the review of this manuscript.

### v3

@@
 So far, no one has clarified how information control emerged ab initio and how primitive control mechanisms in life might have evolved, becoming increasingly refined.
-Based on recent experimental results showing that chemical computation does not require the presence of life-related chemistry (Dueñas-Díez and Pérez-Mercader,, 2019), we elucidate the origin and early evolution of information handling by chemical automata, from information processing (computation) to information storage (memory) and information transmission (communication).
+Based on recent experimental results showing that chemical computation does not require the presence of life-related chemistry (Dueñas-Díez and Pérez-Mercader,, 2019), we elucidate the origin and early evolution of information handling by chemical automata, from information processing (computation) to information storage (memory) and information transmission (communication) and later digital messengers, covering at the same time its syntactic, semantic and pragmatic flavors (Küppers,, 1990).
 In contrast to other theories that assume the existence of initial complex structures, our representation starts from trivial self-replicators whose interaction leads to the arising of more powerful molecular machines.
 By describing precisely the primordial transitions in chemistry-based computation, our framework is capable of explaining the above-mentioned gaps and can be translated to other models of computation, which allow us to explore biological phenomena at multiple spatial and temporal scales.
+Being compatible with the free energy principle, we have developed a computational enactivist (Korbak,, 2021) theoretical framework that could be able to describe from the origin of life to higher-level cognition, as if it were a purely constructivist narrative (Stewart,, 1995).
 At the end of our manuscript, we propose some ways to extend our ideas, including experimental validation of our theory (both in vitro and in silico).
@@
 (Hernández-Orozco et al.,, 2018).
-The undecidability behind self-replication lead us explicitly to open-ended evolution as a result of evolution (Pattee and Sayama,, 2019), and as Hernández-Orozco et al., (2018) has pointed out, systems that exhibit (strong) open-ended evolution must be undecidable.
+The undecidability behind self-replication lead us explicitly to open-ended evolution (OEE) as a result of evolution (Pattee and Sayama,, 2019), and as Hernández-Orozco et al., (2018) has pointed out, systems that exhibit strong OEE must be undecidable.
 The PDA allows the ecosystem to store a more complex set of instructions—including the components of the chemical network that makes up the PDA itself—in a delocalized way.
@@
 Now that we have a “vehicle” in the von Neumann sense, we can talk about self-reproduction (Marletto,, 2015) and proto-organisms (Egbert et al.,, 2023), entities capable of gene-based Darwinian evolution, but possibly lacking some features found in all modern organisms.
+We are converging towards a technological singularity in which the digital is dominant, in such a way that capital is decentralized due to the ease of societies to acquire material goods that allow them to access this digital world.
+In this way, we are heading towards a collective decentralization, in which even material goods will have to be decentralized in order to continue with the socio-techno-economic progress we have seen in the last decades, given by universal indicators such as the global increase in life expectancy, among others.
 Discussion
@@
 Acknowledgements
-The authors acknowledge Luis M. Rocha and three anonymous reviewers for their valuable comments during the review of this manuscript.
+The authors acknowledge Juan Pérez-Mercader, Luis M. Rocha and three anonymous reviewers for their valuable comments during the evaluation of this manuscript.

### v4

@@
 So far, no one has clarified how information control emerged ab initio and how primitive control mechanisms in life might have evolved, becoming increasingly refined.
-Based on recent experimental results showing that chemical computation does not require the presence of life-related chemistry (Dueñas-Díez and Pérez-Mercader,, 2019), we elucidate the origin and early evolution of information handling by chemical automata, from information processing (computation) to information storage (memory) and information transmission (communication) and later digital messengers, covering at the same time its syntactic, semantic and pragmatic flavors (Küppers,, 1990).
+Based on recent experimental results showing that chemical computation does not require the presence of life-related chemistry (Dueñas-Díez and Pérez-Mercader,, 2019), we elucidate the origin and early evolution of information handling by chemical automata, from information processing (computation) to information storage (memory), information transmission (communication) and later digital messengers, covering at the same time their syntactic, semantic and pragmatic flavors (Küppers,, 1990).
 In contrast to other theories that assume the existence of initial complex structures, our representation starts from trivial self-replicators whose interaction leads to the arising of more powerful molecular machines.
 By describing precisely the primordial transitions in chemistry-based computation, our framework is capable of explaining the above-mentioned gaps and can be translated to other models of computation, which allow us to explore biological phenomena at multiple spatial and temporal scales.
-Being compatible with the free energy principle, we have developed a computational enactivist (Korbak,, 2021) theoretical framework that could be able to describe from the origin of life to higher-level cognition, as if it were a purely constructivist narrative (Stewart,, 1995).
+Being compatible with the free energy principle, we have developed a computational enactivist (Korbak,, 2021) theoretical framework that could be able to describe from the origin of life to high-level cognition, as if it were a purely constructivist narrative (Stewart,, 1995).
 At the end of our manuscript, we propose some ways to extend our ideas, including experimental validation of our theory (both in vitro and in silico).
 Introduction
-Leaving aside the philosophical question about the origin of information or the chicken-and-egg problem between information and matter/energy, it is possible to have an even deeper understanding of phenomena such as life or cognition if we assume that information is anything that an entity can sense, perceive, or observe.
-(Gershenson,, 2012).
-Thus, following the line of thought proposed by Pattee, (1969), we can establish the primordial ecosystem as the first information-containing architecture in our planet, whose primitive geochemical cycles gave way to a well-defined syntax—based on chemistry— spontaneously without the need for genetic instructions or initial metabolic control.
-This gave way to molecular Darwinism, a phenomenon that has been well documented for several decades (Küppers,, 1990) and whose formalization was refined the last decade by Marletto, (2015).
-As was explained in Walker and Davies, (2013), neither genetic-first nor metabolism-first approaches can describe completely the emergence of living organisms, as both viewpoints “neglect the active (algorithmic or instructional) and distributed nature of biological information” (Walker and Davies,, 2013).
-In this way, “the real challenge of life’s origin is thus to explain how instructional information control systems emerge naturally and spontaneously from mere molecular dynamics” (Walker and Davies,, 2013).
-Thus, we explore an information-first approach to study the origins and evolution of life.
-Recently, it was shown that chemical reactions can be thought of as molecular recognition machines, which do not need biochemistry to mimic the behavior of recognizing automata belonging to different classes in Chomsky’s hierarchy (Dueñas-Díez and Pérez-Mercader,, 2019).
-The experiments carried out by Dueñas-Díez and Pérez-Mercader, (2019) consisted of implementing strictly “one-pot-reactor” chemical realizations of automata for the recognition of well-known languages at different levels in the Chomsky hierarchy.
+A few years ago it was shown that chemical reactions can be considered molecular recognition machines, which do not need biochemistry to mimic the behavior of recognizing automata belonging to different classes in Chomsky’s hierarchy (Dueñas-Díez and Pérez-Mercader,, 2019).
+These experiments consisted of implementing strictly “one-pot-reactor” chemical realizations of automata.
 To enact a Finite-State Automaton (FSA), they used a simple bimolecular reaction in an aqueous medium.
@@
 To bring about a Turing machine in the Linear Bounded Automaton (LBA) subclass, they used the Belousov-Zhabotinsky (BZ) reaction network.
-In this case, the BZ reaction was able to successfully recognize a context-sensitive language.
+In this case, the BZ reaction was able to successfully recognize a context sensitive language.
 As the authors have pointed out in their conclusions, these experimental results enlighten “a path for the sequential/hierarchical origin of information handling in nature and the transition from non-biochemical to biochemical and extant life-related computation” (Dueñas-Díez and Pérez-Mercader,, 2019).
-There are several theoretical frameworks which describe life as information (Chaitin,, 2012; Farnsworth et al.,, 2013; Marletto,, 2015), but none of them has explained how information handling emerged from chemical interactions.
+Today there are several theoretical frameworks which describe life as information (Chaitin,, 2012; Farnsworth et al.,, 2013; Marletto,, 2015), but none of them has explained how information handling emerged from chemical interactions.
 By information handling, we mean information processing (computation), information storage (memory), and information transmission (communication).
 As will be discussed below, such information handling also includes the syntactic, semantic and pragmatic flavors of information.
-In this work, we describe precisely the primordial transitions in chemistry-based computation, from the appearance of trivial replicators (simple bimolecular reactions, equivalent to FSAs) in the environment up to the evolution of approximated chemical Turing machines (complex chemical networks, represented here by LBAs) and, eventually, to their combination into more powerful molecular computers, capable of performing Bayesian inference (Poole et al.,, 2017) and associative learning (Bartlett and Louapre,, 2022), which could have led to the development of autonomy (Kirchhoff et al.,, 2018) and high-level cognition (Levin,, 2019).
+As was explained in Walker and Davies, (2013), neither genetic-first nor metabolism-first approaches can describe completely the emergence of living organisms, as both viewpoints “neglect the active (algorithmic or instructional) and distributed nature of biological information” (Walker and Davies,, 2013).
+In this way, “the real challenge of life’s origin is thus to explain how instructional information control systems emerge naturally and spontaneously from mere molecular dynamics” (Walker and Davies,, 2013).
+Thus, we will explore an information-first approach to study the origins and evolution of life.
+Leaving aside the philosophical question about the origin of information or the chicken-and-egg problem between information and matter/energy, it is possible to have an even deeper understanding of phenomena such as life or cognition if we assume that information is anything that an entity can sense, perceive, or observe (Gershenson,, 2012).
+Thus, following the line of thought proposed by Pattee, (1969), we can establish the primordial ecosystem as the first information-containing architecture in our planet, whose primitive geochemical cycles gave way to a well-defined syntax—based on chemistry—spontaneously without the need for genetic instructions or initial metabolic control.
+This gave way to molecular Darwinism, a phenomenon that has been well documented (Küppers,, 1990) and whose formalization was refined the last decade by Marletto, (2015).
+In this work, we describe precisely the primordial transitions in chemistry-based computation, from the appearance of trivial replicators—simple bimolecular reactions, equivalent to FSAs—in the environment up to the evolution of approximated chemical Turing machines—complex chemical networks, represented here by LBAs—and, eventually, to their combination into more powerful molecular computers, capable of performing Bayesian inference (Poole et al.,, 2017) and associative learning (Bartlett and Louapre,, 2022), which could have led to the development of autonomy (Kirchhoff et al.,, 2018) and high-level cognition (Levin,, 2019).
 For us, the presence of FSAs in the environment demarcates the origin of information processing (computation), PDAs define the appearance of information storage (memory) by the environment and LBAs bound the emergence of information transmission (communication) through different entities in the environment.
@@
 Information Processing
-As has been suggested by Chaitin, (2012), think of our planet as a big parallel, quasi-universal Turing machine capable of simulating a huge variety of functions (programs) can give us novel insights about evolution.
+As has been suggested by Chaitin, (2012), thinking of our planet as a big parallel, quasi-universal Turing machine capable of simulating a huge variety of programs can give us novel insights about evolution.
 Here, the term “quasi-universal” is due to the fact that “real” universal Turing machines are pure abstractions which need an infinite memory tape.
 Natural resources are finite, and this is a first limitation for life to emerge.
-Extrapolating the ideas suggested in Hernández-Orozco et al., (2018) and inspired by the evolutionary model proposed in von Neumann, (1966), the first organisms on the planet could be described as some kind of primitive automata which were assembled and executed by these primeval habitats.
-As noted above, the primordial geophysical and geochemical constraints generated the primeval ecosystem language (Pattee,, 1969), which was interpreted by such primitive entities.
-The cyclic nature of such processes allows for a robust flow of information—abolishing the fragility or deterioration of it—, resulting in the emergence of elementary autocatalytic bimolecular reactions.
-These chemical reactions are capable of recognizing regular languages, a task that does not involve counting or memory (Dueñas-Díez and Pérez-Mercader,, 2019).
-By the good regulator theorem (Conant and Ross Ashby,, 1970), some of the elementary bimolecular reactions in that primeval environment must have a cyclic nature similar to that of geologic phenomena.
-Thus, it is possible to achieve self-replication by simple bimolecular reactions.
-Since it involves the unassisted formation of a covalent bond between the two reactants, the reactions will be relatively slow (Duim and Otto,, 2017).
-Nevertheless, the autocatalytic cycle will play an increasingly dominant role (Kauffman,, 1986) if a sufficiently large amount of the template molecules is formed (Duim and Otto,, 2017), which in part could also promote self-replication.
-Self-replication differs from self-reproduction in the existence of a copier $C$ acting on raw material (Szathmáry,, 1995).
+Inspired by von Neumann, (1966) and extrapolating the ideas suggested in Hernández-Orozco et al., (2018), we propose that the first organisms on the planet are described as primitive automata which were assembled and executed by these primeval habitats.
+Recently, it was shown that “when random, non self-replicating programs are placed in an environment lacking any explicit fitness landscape, self-replicators tend to arise” (Alakuijala et al.,, 2024).
+Thus, given a set of primordial geophysical and geochemical constraints—a primeval ecosystem language (Pattee,, 1969)—we can glimpse the appearance of simple bimolecular reactions that later became autocatalytic, being capable of self-replication.
+Here is important to note that self-replication differs from self-reproduction in the existence of a copier $C$ acting on raw material (Szathmáry,, 1995).
 As has been well argued by Marletto, (2015): “for looser replicators, such as crystals, or short RNA strands and autocatalytic cycles involved in the origin of life, the copier $C$ is null, i.e.
@@
 Although it has been speculated that the simple interaction between self-replicators and the environment leads to open-ended evolution, for us this is not a prerequisite.
-These trivial replicators, as Walker and Davies, (2013) call them, are simple enough to have been naturally occurring in pre-biological environments and kick off the origin of information processing.
-These structures do not require a memory nor a “vehicle”—in the von Neumann sense (von Neumann,, 1966)—, so we can imagine the architecture of these earliest chemical machines as being represented by autocatalytic bimolecular chemical reactions, which under our framework are single FSAs (Dueñas-Díez and Pérez-Mercader,, 2019), capable of constructing themselves using natural resources and a simple description encoded in the environment.
+These trivial replicators, as Walker and Davies, (2013) call them, are capable of recognizing regular languages, a task that does not involve counting or memory (Dueñas-Díez and Pérez-Mercader,, 2019), kicking off the origin of information processing.
+In addition, they do not require a “vehicle”—in the von Neumann sense (von Neumann,, 1966)—so we can imagine the architecture of these earliest chemical machines as being represented by autocatalytic bimolecular chemical reactions, which under our framework are single FSAs (Dueñas-Díez and Pérez-Mercader,, 2019), capable of constructing themselves using natural resources and a simple description encoded in the environment.
 Just as $L$-systems, it is clear that the algorithmic information of these simple architectures is one that may be specified by an algorithm containing far fewer bits than the structure it describes.
-At this point, it is important to note that the different dimensions of the concept of information (syntactic, semantic and pragmatic) cannot be separated from each other (Küppers,, 1990).
+A general concern up to this point is the fragility of information.
+How can we assert that information will be conserved through these autocatalytic bimolecular reactions?
+First of all, even in a primordial stage, the nature of the geophysical-chemical processes is cyclical.
+This fact ensures that there is a constant energized, non-equilibrium, dynamic, cyclic exchange between the environment and the self-replicators.
+Under this condition we can talk about dynamic kinetic stability (DKS), a kinetic form of stability “associated with entities able to make copies of themselves at a rate that results in a non-equilibrium steady-state population of replicating entities being maintained over time—persistence through self-replication” (Pascal and Pross,, 2015).
+Given that DKS guides self-replication towards a more robust replication state (Pascal et al.,, 2023) and increasingly complex dynamics continue to emerge following the rise of self-replicators (Alakuijala et al.,, 2024), there must be a preservation of the encoded information in the environment, giving way to molecular Darwinism and at the same time abolishing the deterioration of information.
+At this point, it is important to note that the different dimensions of the concept of information—syntactic, semantic and pragmatic—cannot be separated from each other (Küppers,, 1990).
 Although the existence of semantic (or pragmatic) information in such a primordial state of life may seem far-fetched, it must first be remembered that meaning is a multi-scale phenomenon (Fields and Levin,, 2020).
@@
 Information Storage
-So far, we have a world in which bimolecular reactions (represented here by FSAs) coexist.
+So far, we have a world in which bimolecular reactions—represented here by FSAs—coexist.
 These are capable of computation and self-replication but have no vehicles or memory.
@@
 Under this panorama, some chemical intermediates are acting as the product of one reaction and as reactant of another one, forming networks of even more complex chemical reactions.
-This dynamical process is identical to the premise of autocatalysis and is essentially equivalent to the push/pop duality in one-stack PDAs described in (Dueñas-Díez and Pérez-Mercader,, 2019).
-The transition from FSAs (simple bimolecular reactions) to one-stack PDAs (networks of chemical reactions) demarcates the appearance of a primitive information storage (memory) by the environment, because it allows for the recognition of more complex sequences of information (Hopcroft et al.,, 2001) and thus it empowers a higher evolutionary robustness of such information in the face of changing environmental conditions (Walker and Davies,, 2013).
+This dynamical process is identical to the premise of autocatalysis and is essentially equivalent to the push/pop duality in one-stack PDAs described in Dueñas-Díez and Pérez-Mercader, (2019).
+Note that the transition from FSAs (simple bimolecular reactions) to one-stack PDAs (networks of chemical reactions) demarcates the appearance of a primitive information storage (memory) by the environment, because it allows for the recognition of more complex sequences of information (Hopcroft et al.,, 2001) and thus it empowers a higher evolutionary robustness of such information in the face of changing environmental conditions (Walker and Davies,, 2013).
+A widespread misconception in literature is the idea that computation and construction are two different phenomena.
+We argue that this is due to the fact that in his seminal work (von Neumann,, 1966), John von Neumann did not clarify the relationship between his universal constructor and the universal Turing machine proposed decades earlier by Turing, (1936).
+Nevertheless, by Kleene’s second recursion theorem we know that Turing machines can obtain their own description and then go on to compute with it (Sipser,, 1996).
+Since the recursion theorem shows how machines can reproduce themselves, both computation and construction are two sides of the same coin.
+In Sayama, (2008) one of us declared that the difference between the replicating programs of recursion theory and the self-replicating constructors in (von Neumann,, 1966) is that the latter never halt.
+However, this is not stated anywhere in recursion theory (Sipser,, 1996).
+In fact, one application of the recursion theorem is quines, computer programs that receive no input and produce a copy of their own source code as their only output.
+If we run a quine in an infinite loop, it will never halt, and since the recursion theorem is independent of the computational substrate (chemical reactions in our case), we claim that construction is a corollary of recursion theory.
+Once we have explained the above, it is important to note that in this context undecidability is still synonymous with self-replication (Sayama,, 2008), since it is in this indeterminate state where we will be certain that the construction process will continue.
+Given that the environment—described here as a quasi-universal computer compartment—“is constantly approximating solutions to the adaptation problem by means of survivability” (Hernández-Orozco et al.,, 2018), nature pursuits of the busy beaver values (longest-running automata) in the environment, which is identical to the search for the longest-surviving organisms.
+At the level of the automata (a chemical network), when these attempt to execute the instruction of being the longest-running machine, it produces a nested loop of computation which in terms of construction translates as self-replication (Sayama,, 2008).
+This process is summarized in Fig.
+2.
 From a von Neumann mindset, the one-stack PDA is now some kind of primitive constructor, because we have a chemical network that might be able to reproduce itself including the components that conform it.
-At this computational level, self-replication can be seen as the result of nature trying to solve a equivalent PDA-problem to the Halting problem (Sayama,, 2008).
-Following our computational train of thought, the ecosystem, a quasi-universal computer compartment, “is constantly approximating solutions to the adaptation problem by means of survivability” (Hernández-Orozco et al.,, 2018).
-In this way, nature pursuits of the busy beaver values (longest-running automata) in the environment, which is identical to the search for the longest-surviving organisms.
 At this particular evolutionary stage, nature is looking for which PDA will be the longest running machine, which is interchangeable to find the most robust primitive information storage for the environment.
 This is equivalent to computing the largest number of steps taken by an $n$-state, $m$-symbol machine started on an initially non-blank tape before halting.
-This is of course an undecidable problem comparable to a busy beaver game (Rado,, 1962), and at the level of the PDA, when this attempts to execute the instruction of being the longest-running machine, it produces a nested loop of computation which in terms of construction translates as self-replication (Sayama,, 2008).
-This process is summarized in Fig.
-2.
+This is of course an undecidable problem comparable to a busy beaver game (Rado,, 1962).
+Thus, up to this computational level, self-replication can be seen as the result of nature trying to solve a PDA-problem equivalent to the Halting problem (Sayama,, 2008).
 Figure 2: Explanation of PDA and LBA self-replication.
-First, the environment (described as a quasi-universal computer compartment) approximates solutions to the adaptation problem by means of survivability, which guides the search for the longest-surviving organisms.
+First, the environment—described as a quasi-universal computer compartment—approximates solutions to the adaptation problem by means of survivability, which guides the search for the longest-surviving organisms.
 Next, the automata (a chemical network) attempt to execute the instruction of being the longest-running machine.
@@
 (Hernández-Orozco et al.,, 2018).
-The undecidability behind self-replication lead us explicitly to open-ended evolution (OEE) as a result of evolution (Pattee and Sayama,, 2019), and as Hernández-Orozco et al., (2018) has pointed out, systems that exhibit strong OEE must be undecidable.
+The undecidability behind self-replication lead us explicitly to open-ended evolution (OEE) as an output of evolution (Pattee and Sayama,, 2019), and as Hernández-Orozco et al., (2018) has pointed out, systems that exhibit strong OEE must be undecidable.
 The PDA allows the ecosystem to store a more complex set of instructions—including the components of the chemical network that makes up the PDA itself—in a delocalized way.
@@
 At this stage we can talk about ante-organisms, life-like systems that are capable of self-preserving behaviour (Egbert et al.,, 2023).
-Up to this point we have a resource-limited ecosystem inhabited by chemical networks that compete, self-replicate and mutate (through small variations in their composition) in such a way that each kind of replicator produce at least one viable offspring, on average, per lifetime.
+Up to this point we have a resource-limited ecosystem inhabited by chemical networks that compete, self-replicate and mutate—through small variations in their composition—in such a way that each kind of replicator produce at least one viable offspring, on average, per lifetime.
 Therefore, we can claim that so far, our ante-organisms are capable of primitive evolution (Marletto,, 2015), which is a kind of natural selection with limited inheritance caused by low robustness and accuracy in replicating the information stored in the chemical network by the environment.
@@
 The first one is through the complexification of one initial PDA.
-Since at this stage the entities are able to propagate mutations across their offspring (represented by variations in the structures of the chemical networks), it is possible that they can become more and more complex, in a way that they can develop two or more push/pop dualities.
+Since at this stage the entities are able to propagate mutations across their offspring—represented by variations in the structures of the chemical networks—, it is possible that they can become more and more complex, in a way that they can develop two or more push/pop dualities.
 The second way is essentially through the coupling of two or more PDAs.
@@
 In both cases we obtain even more complex chemical networks that allow the identification of more abstract sequences of information, resulting in an even more sophisticated information storage system on the part of the environment.
-Because chemical automata need communication in some way to cooperate, this era launches information transmission.
+Because chemical automata need to communicate in some way in order to cooperate, this era sets in motion the transmission of information.
 Given a set of selfish agents minimizing their stress (surprise) and competing for information, it is possible to develop cooperative homeostatic loops which improve the predictive capabilities of the entities and gives rise to functional relationships (Levin,, 2019; Heins et al.,, 2024).
@@
 Notice that at this stage the abstract and non-physical systemic entities effectively become causal agents capable of manipulating more and more of its material substrate (Auletta et al.,, 2008).
-Following the mindset of (Walker and Davies,, 2013), we are having a transition from trivial to non-trivial replicators, where the latter requires an algorithm, or instruction set, of complexity comparable to the system it describes (or creates).
+Following the mindset of Walker and Davies, (2013), we are having a transition from trivial to non-trivial replicators, where the latter requires an algorithm, or instruction set, of complexity comparable to the system it describes (or creates).
 In the same way, we can imagine this metamorphosis from reactive to representational systems (Zenil et al.,, 2012).
@@
 Certainly, this process could have taken a long time to happen, such that digital life forms may be “the only systems that survive in the long-run and are thus the only remaining product of the processes that led to life” (Walker and Davies,, 2013).
-The appearance of digital messengers can be seen as the implementation of descriptions that can (or cannot) interact with the primitive chemical constructors to be constructed or copied.
+The appearance of digital messengers could be seen as the implementation of descriptions that can (or cannot) interact with the primitive chemical constructors to be constructed or copied.
 Of course, this is not a necessary condition for life, and it will depend on which architecture we are talking about.
@@
 Now that we have a “vehicle” in the von Neumann sense, we can talk about self-reproduction (Marletto,, 2015) and proto-organisms (Egbert et al.,, 2023), entities capable of gene-based Darwinian evolution, but possibly lacking some features found in all modern organisms.
-We are converging towards a technological singularity in which the digital is dominant, in such a way that capital is decentralized due to the ease of societies to acquire material goods that allow them to access this digital world.
-In this way, we are heading towards a collective decentralization, in which even material goods will have to be decentralized in order to continue with the socio-techno-economic progress we have seen in the last decades, given by universal indicators such as the global increase in life expectancy, among others.
 Discussion
@@
 The real challenge of the origin of life is not the existence of such chemical computation, but reliable computation capable of permeating the protobiosphere.
-Particularly, the framework proposed by Chollet, (2019) seems suitable to quantify how resilient the chemical entities proposed here are in the face of a changing environment.
+Particularly, an extension of the framework proposed by Chollet, (2019) seems suitable to quantify how resilient the chemical entities proposed here are in the face of a changing environment.
 In order for there to be such automata capable of interpreting and storing the information contained in the environment, there must first be a primordial language (conformed by chemicals) in the environment that allows the realization of such automata.
@@
 We have used many computational concepts when describing our ideas, however, our theory points to the fact that life and cognition are two intertwined, parallel phenomena.
-As has been shown in Stewart, (1995), it is possible to describe from life’s origin to higher-level cognition using purely a constructivist description.
+As has been shown in Stewart, (1995), it is possible to describe from life’s origin to high-level cognition using purely a constructivist description.
 Being compatible with the free energy principle, in this paper we have developed a computational enactivist (Korbak,, 2021) theoretical framework capable of being as explanatory as a strictly constructivist approach.
@@
 Since our theory is compatible with the von Neumann’s replicator-vehicle logic (von Neumann,, 1966), it is possible under no-design laws (Marletto,, 2015), provided that the latter permit information media and enough generic resources.
-In this way, we have already proposed the path for the sequential/hierarchical origin of information handling in nature and the transition from non-biochemical to biochemical and extant life-related computation.
+In this way, we just have proposed the path for the sequential/hierarchical origin of information handling in nature and the transition from non-biochemical to biochemical and extant life-related computation.
 Notwithstanding, even though our narrative describes how the primitive information control mechanisms in life might have evolved, we do not claim from what point life has emerged.
@@
 Given that under our perspective life is describable at all extents, there should not be privileged level of causation (Noble,, 2012).
-This opens the possibility to characterize phenomena like downward causation (Flack,, 2017) from our theoretical framework, and even if the informational-causal narrative is accepted as a mere façon de parler, our metaphor is still compatible with theories where only material objects enjoy true causal efficacy (Walker et al.,, 2012).
+This opens the possibility to characterize phenomena like downward causation (Flack,, 2017) from our theoretical framework, and even if the informational-causal narrative is accepted as a mere façon de parler, our framework is still compatible with theories where only material objects enjoy true causal efficacy (Walker et al.,, 2012).
 Future work
@@
 Of the many approaches that exist in synthetic biology (Hanczyc,, 2020), protocells offers a comprehensive resource on current attempts to create simple forms of life from scratch in the laboratory (Deamer,, 2008).
-In fact, using the BZ reaction (equivalent to a LBA) it is possible to devise active synthetic polymer vesicles, which self-reproduce (Bastakoti and Perez-Mercader,, 2017) and evolve (Katla et al.,, 2023).
+In fact, using the BZ reaction (equivalent to a LBA) it is possible to devise active synthetic polymer vesicles, which self-reproduce (Bastakoti and Perez-Mercader,, 2017) and evolve primitively (Katla et al.,, 2023).
 Is it possible to synthesize vesicles using a pH chemical reaction (equivalent to a PDA) or a precipitation reaction (equivalent to a FSA)?
@@
 The same research group has obtained three reaction-diffusion equations that are capable of representing information processing, metabolism, self-replication and evolution (Muñuzuri and Pérez-Mercader,, 2022).
-The solutions to these equations show the phenomenology of a generic living system.
+In fact, the solutions to these equations show the phenomenology of a generic living system.
 To obtain these equations the authors implemented each of the four properties (information processing, metabolism, self-replication and evolution) one by one.
@@
 We can use as input the chemical components and reactions proposed in Dueñas-Díez and Pérez-Mercader, (2019) for every class of chemical automata and observe which types of minimal paths of organization are formed.
-This is something similar to the methodology proposed by Fontana and Buss, (1994), which was refined later by Szathmáry, (1995).
+This is something similar to the methodology proposed by Fontana and Buss, (1994), which was refined later by Szathmáry, (1995) and recently inspired the work by Alakuijala et al., (2024).
 Are these organizations able to self-reproduce, compute, have memory, or communicate?
@@
 Together with our description of how information handling appeared, viability-based behaviour it is a good candidate to explore how agency has emerged and evolved, or to model scale-free cognition (Levin,, 2019).
-In this way, instead of creating new theoretical frameworks and computational tools to study phenomena such as agency, life, intelligence, identity, and cognition, we promote the unification of what already exists by looking for global equivalences among different formalisms.
+In this way, instead of creating new theoretical frameworks and computational tools to study phenomena such as agency, life, intelligence, identity, and cognition, we promote the unification of what already exists by looking for global equivalences among different theories.
 Types of organisms Corresponding automata
@@
 Acknowledgements
-The authors acknowledge Juan Pérez-Mercader, Luis M. Rocha and three anonymous reviewers for their valuable comments during the evaluation of this manuscript.
+The authors acknowledge Juan Pérez-Mercader, Luis M. Rocha, David Wolpert and three anonymous reviewers for their valuable comments.